What two sets of reagents (each consisting of a carbonylcompound and phosphonium ylide) can be...

  1. What two sets of reagents (each consisting of a carbonyl compound and phosphonium ylide) can be used for the synthesis of each of the following alkenes?

  2. What alkyl halide is required to prepare each of the phosphonium ylides?

  3. What is the best set of reagents to use for the synthesis?
Answer #1

2 (C6H5) C = CH CH₂ Two sets of Reagents are 1 (CHS) c=0 + CH₃ CH = ppha Carbonys compound. Ylide 2) (CHS) C = pphz + CH₃-CHO